### v1

# QCSE: A Pretrained Quantum Context-Sensitive Word Embedding for Natural Language Processing

Abstract
Quantum Natural Language Processing (QNLP) offers a novel approach to encoding and understanding the complexity of natural languages through the power of quantum computation. This paper presents a pretrained quantum context-sensitive embedding model, called QCSE, that captures context-sensitive word embeddings, leveraging the unique properties of quantum systems to learn contextual relationships in languages. The model introduces quantum native context learning, enabling the utilization of quantum computers for linguistic tasks. Central to the proposed approach are innovative context matrix computation methods, designed to create embeddings of words based on their surrounding linguistic context. Five distinct methods are proposed and tested for computing the context matrices, incorporating techniques such as exponential decay, sinusoidal modulation, phase shifts, and hash-based transformations. These methods ensure that the quantum embeddings retain context sensitivity, thereby making them suitable for downstream language tasks where the expressibility and properties of quantum systems are valuable resources. To evaluate the effectiveness of the model and the associated context matrix methods, evaluations are conducted on both a Fulani corpus, a low-resource African language, a small-size dataset, and an English corpus of slightly larger size. The results demonstrate that QCSE not only captures context sensitivity, but also leverages the expressibility of quantum systems to represent rich, context-aware language information. The use of Fulani further highlights the potential of QNLP to mitigate the problem of lack of data for this category of languages. This work underscores the power of quantum computation in natural language processing (NLP) and opens new avenues for applying QNLP to real-world linguistic challenges across various tasks and domains.
Index Terms:
Quantum Natural Language Processing, Quantum Word Sequence Embedding, Natural Language Processing, Machine learning, Quantum Computing, Quantum Machine Learning, Word Embedding.
[…]
=-21pt
I Introduction
Natural Language Processing (NLP) has rapidly become a cornerstone of modern technology and industry, playing a key role in applications ranging from language translation and text generation, to virtual assistants and data understanding and analysis [1, 2, 3, 4]. These advances have been driven by powerful machine learning models that enable computers to process, understand, reason, and generate human language [5, 6, 7]. Despite this progress, NLP models still face significant challenges, especially when it comes to understanding complex context, meaning, and nuances in language[8, 9]. Classical models are computationally expensive and still face limitations in capturing deep contextual meaning, struggling with certain subtleties in natural language, especially in complex language structures[10, 11, 12, 8, 9, 13, 14, 15, 16, 17, 18]. As the demand for more efficient and accurate NLP solutions grows, researchers are looking for new approaches to push the boundaries. Quantum computing [19, 20, 21, 22] offers a promising frontier in this regard.
With its ability to process information in ways that classical computers cannot, quantum computing opens up new computation possibilities. Quantum machine learning (QML) [23, 24, 25, 26] and QNLP [27, 22, 28] explore how the principles of quantum mechanics — like superposition and entanglement — can be applied to understand language more effectively. In recent years, several pioneering works have laid the foundations for QML in NLP. Following works [29, 30, 25, 31] provide a foundational understanding of QML, explaining how quantum computers could enable new learning algorithms that go beyond classical machine learning. Although QNLP is still in its infancy, its potential to surpass classical embeddings is being actively investigated, potentially leading to breakthroughs in how language is modeled and understood. Pioneering works on QNLP models, such as the quantum language model (QLM) [32, 33] has been explored for a variety of NLP tasks, including question answering and sentiment analysis[34, 35, 36], where superposition and entanglement-based embeddings have shown promising results in capturing dependencies between words in sentences. The rise of QNLP has sparked interest in developing particularly quantum word and sequence embedding models [37, 38, 39, 22, 40].
Quantum word embedding aims to represent words as quantum states that inherently capture their context and meaning for subsequent NLP tasks. A wide range of quantum embedding models has been proposed to handle specific language representation challenges. One such model is the complex-valued embedding model (CVE) [37, 41, 42, 43], which uses quantum-inspired representations to map words onto complex-valued vectors, capturing richer information compared to real-valued embeddings. The CVE model has been applied with promising results to NLP tasks such as sentiment analysis and classification [44, 45, 46]. Similarly, the quantum-inspired pretrained model [47] is designed to integrate quantum principles with classical deep neural networks, offering a hybrid approach to learning word embeddings with pretraining strategies commonly found in NLP. Other recent contributions, which exploit quantum entanglement are word2ket [38] and the quantum language model with entanglement embedding (QLM-EE) [32]. QLM-EE adapts classical pretrained architectures with quantum principles for text understanding. These models mark significant progress in the introduction of quantum-based techniques to NLP and open the door to future innovations in the field.
Despite their potential, current quantum embedding models face limitations that hinder their ability to fully exploit the advantages of quantum computing. One of the most notable limitations is their reliance on pre-trained classical embedding models, such as GloVe, RoBERTa, or word2vec[44, 32, 47] or random initialization, to generate initial word vectors. These learned classical embeddings or vectors are then mapped onto quantum states, rather than training the quantum embedding models directly from the corpus to learn quantum-native and contextual relationships. This dependence could lead to the propagation of the inherent limitations of classical models through the subsequent stages of the quantum process by inheriting the constraints of classical representations.
In this work, we design a quantum word embedding model for learning quantum embeddings directly from the corpus, without relying on pretrained classical embeddings or random initialization. By learning initial word contexts in a purely quantum-native framework, the model could better capture the polysemy and complex semantic relationships that are inherent in natural language.
The purpose of this work is to explore the potential of quantum word and sequence embedding and to discuss how it could enhance the performance and capabilities of NLP systems. The key contributions are the following.
  • Introduce a quantum embedding model that directly learns the context of words from the corpus, bypassing the need for pre-trained classical word embeddings[44, 32, 47]. Unlike existing models that rely on classical pre-trained word vectors, the proposed quantum model directly learns from the word co-occurrence data in the corpus. This approach enables quantum states to be constructed without dependency on traditional embedding models, offering a fully quantum-native representation of words and their contextual relationships.
  • Various methods are proposed for computing context matrices that serve as the foundation for encoding word relationships into quantum states. These techniques ensure that word proximity and dependencies are effectively captured within quantum circuits, facilitating dynamic quantum word embedding.
  • Propose a quantum circuit with the balance between circuit depth and expressibility. The quantum circuits are optimized to efficiently encode word contexts with a manageable and constant number of qubits, ensuring scalability of the model for larger vocabularies and longer sequences.
  • In the end, the paper provides an evaluation of the model’s performance across different dataset sizes. The losses and precision of quantum embeddings are evaluated, illustrating the advantages of quantum-native approaches in terms of context capture and semantic precision.
The remainder of this paper is organized as follows. Section II provides an overview of NLP and QNLP, focusing on key concepts for understanding sake and transitioning to QNLP. In particular, Section II-A introduces the foundational principles of quantum computing concepts, essential for understanding the subsequent sections. Section II-B explores the background of QNLP, with a focus on quantum information encoding. Section III details the proposed quantum word embedding model, including the model overview and its different components. Section III-D provides a complexity analysis of the model in terms of number of qubits and quantum gates. Section IV discusses the different methods for computing context matrices and vectors, which are critical to capturing and encoding semantic relationships in quantum circuits. To asses the performance of the model, Section V presents the evaluations carried out to validate our approach, followed by a discussion of the results. From the discussion, Section VI outlines future directions for research in QNLP, and finally Section VII concludes the paper with a summary of our findings and their implications.
II Preliminaries
II-A Quantum Computing Concepts
To provide a foundational understanding of the key concepts driving quantum computing and QNLP, this section introduces the fundamental principles of bits vs. qubits, superposition, entanglement, and quantum states with $m$ qubits, along with their mathematical formulations [48]. These concepts form the basis for exploring quantum word embeddings and their applications in NLP.
Definition 1
A classical bit is the fundamental unit of classical information, which can exist in one of two discrete states, 0 or $1$. A qubit is the quantum analog of a classical bit. Unlike a classical bit, a qubit can exist in a superposition of states, represented as:
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ (1)
where $\alpha,\beta\in\mathbb{C}$ are complex numbers satisfying the normalization condition:
$|\alpha|^{2}+|\beta|^{2}=1.$ (2)
Here, $|0\rangle$ and $|1\rangle$ are the basis states of the qubit.
Definition 2
Superposition is a fundamental principle of quantum mechanics that allows a quantum system to exist in multiple states simultaneously. For a single qubit, the state is a linear combination of the basis states, as in Equation (1).
Thus, a quantum state with $m$ qubits is described by a vector in a $2^{m}$-dimensional Hilbert space. The general form of an $m$-qubit quantum state is
$|\psi\rangle=\sum_{i=0}^{2^{m}-1}\alpha_{i}|i\rangle,$ (3)
where $|i\rangle$ represents the $i$-th computational basis state (e.g., $|000\ldots 0\rangle,|000\ldots 1\rangle,\ldots,|111\ldots 1\rangle$), and $\alpha_{i}$ are complex amplitudes satisfying the normalization condition:
$\sum_{i=0}^{2^{m}-1}|\alpha_{i}|^{2}=1$ (4)
Definition 3
Entanglement[49, 50] is a quantum phenomenon in which two or more quantum systems become correlated in such a way that the state of one cannot be described independently of the others. The most simple relevant entangled states are the Bell states. One of these Bell states can be written as
$|\psi\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle).$ (5)
This state cannot be factored or separated into individual qubit states, i.e., $|\psi\rangle\notin\{|\psi_{1}\rangle\otimes|\psi_{2}\rangle:|\psi_{1}\rangle,|\psi_{2}\rangle\in\mathcal{H}\}$.
Definition 4
Quantum gates are reversible operations that manipulate the states of qubits, enabling superposition and entanglement to process quantum information. Single-qubit Pauli gates act as
$X|0\rangle=|1\rangle,\quad X|1\rangle=|0\rangle\quad\text{(bit flip)},$
$Z|0\rangle=|0\rangle,\quad Z|1\rangle=-|1\rangle\quad\text{(phase flip)},$
$Y|0\rangle=i|1\rangle,\quad Y|1\rangle=-i|0\rangle\quad\text{(bit and phase flip)},$
with matrix representations:
$X$ $=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. (6)
On a composite Hilbert space $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$, two-qubit gates include the controlled-NOT that flips the target qubit if the control qubit is $|1\rangle$:
$\text{CNOT}_{1,2}=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes X.$ (7)
The CNOT gate can be used to generate Bell states:
$\text{CNOT}_{1,2}\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle)\otimes|0\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle).$
The controlled-Z gate applies a phase flip when both qubits are $|1\rangle$:
$\text{CZ}_{1,2}=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes Z=\text{diag}(1,1,1,-1).$ (8)
II-B Quantum Natural Language Processing
Natural Language Processing involves two key stages: word embedding and model training. Given a vocabulary $V$, words are mapped to $d$-dimensional vectors via an embedding function $f:V\rightarrow\mathbb{R}^{d}$. Popular methods include word2vec [51] and GloVe [52] for maximizing the log-likelihood of context words $w_{c}$ given center word $w$:
$\mathcal{L}=\sum_{(w,w_{c})\in D}\log P(w_{c}|w),\quad P(w_{c}|w)=\text{softmax}(Ww_{c}+b)$ (9)
where $W\in\mathbb{R}^{k\times d}$ and $b\in\mathbb{R}^{k}$ are learnable parameters.
These embeddings $x_{1},...,x_{T}$} are then processed by sequence models such as RNNs/LSTMs [53, 54, 55] and transformers [56, 57] for learning patterns over sequences of words. In a recurrent model like LSTM, the hidden states $h_{t}$ are computed as a function of the previous hidden states and the current input: $h_{t}=f(h_{t-1},x_{t})$. $h_{t}=f(h_{t-1},x_{t})$. The training minimizes for example a cross-entropy loss $\mathcal{L}_{\text{train}}=-\sum_{i=1}^{N}y_{i}\log\hat{y}_{i}$, where $y_{i}$ is the true label and $\hat{y}_{i}$ is the predicted probability for sample $i$.
QNLP leverages the principles of quantum computing to enhance traditional NLP tasks. A key concept in QNLP is the use of quantum word embedding, where classical data (words or sentences) are embedded into quantum states. Instead of representing words classically in a high-dimensional vector space, each word is encoded in a quantum state $|\psi\rangle$ in a Hilbert space. Mathematically, this can be seen as mapping each word $w$ in the vocabulary $V$ to a quantum state $|\psi_{w}\rangle$, where the amplitude or angle of each qubit encodes information about the context of the word. For instance, using amplitude encoding, a vector $\mathbf{v}\in\mathbb{R}^{d}$ representing a word can be normalized and encoded into a quantum state $|\psi\rangle=\sum_{i=1}^{d}v_{i}|i\rangle$. Quantum word embeddings could leverage the superposition and entanglement properties of quantum systems to capture richer contextual relationships between words than classical embeddings.
Training QNLP models typically involves a QML framework such as the variational quantum circuit (VQC)[58, 29, 59], where the circuit parameters are optimized to minimize a loss function. Let $|\psi_{w}\rangle$ be the initial quantum state representing a word or sentence. The VQC transforms this state into another state through a series of unitary operations, $|\psi_{\text{output}}\rangle=U(\theta)|\psi_{w}\rangle$, where $U(\theta)$ is a parameterized unitary matrix dependent on the parameters $\theta$. These parameters are updated via classical optimization algorithms, such as gradient descent, minimizing a loss function $\mathcal{L}(\theta)$ computed from quantum measurements. In QNLP, such training often involves measuring the expectation values of quantum operators to compute probabilities for different outcomes, similar to classical model training. In the case of the quantum embedding model, the final model aims at maximizing the likelihood $P(|w_{\text{context}}\rangle||w_{\text{center}\rangle},\theta)$, where $|w_{\text{center}}\rangle$ and $|w_{\text{context}}\rangle$ are quantum states of the word to embed and its context words in a sentence, respectively, similar to the classical word embedding prediction tasks. This enables the model to learn contextual word embeddings in a quantum-enhanced space. Quantum encoding strategies are crucial for effectively transforming text into processable quantum states, offering potential advantages over classical methods in contextual representation.
III Proposed Quantum Word Embedding Model
III-A Model overview
The aim is to map each word to a quantum state that captures its semantic meaning in the context of a given task. Suppose that we have a set of words $W=\{w_{1},w_{2},\dots,w_{n}\}$, and a vocabulary of size $|V|$. In classical word embeddings, each word is mapped to a high-dimensional vector $\mathbf{w}_{i}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. For quantum embeddings, the quantum model instead maps words to quantum states $|w_{i}\rangle\in\mathcal{H}_{m}$, where $\mathcal{H}_{m}$ is a Hilbert space of dimension $m$, and the quantum embedding process is represented by a unitary transformation that incorporates the word context. The process of learning quantum embeddings involves applying a VQC, typically referred to as an ansatz, to capture the relationships between words in the context of a given task. The ansatz parameters are updated iteratively using optimization methods, such as gradient-based optimization, to minimize a task-specific loss function, such as the cross-entropy. For a quantum sequence embedding model, words in a sentence are encoded as sequences of quantum states, and entanglement operations can capture the dependencies between words in a sequence. We propose a context matrix concept (see Section IV) that transforms the surrounding terms of a word into a structured matrix that captures identifiers, positions, distances, and semantic–syntactic relations to construct the context-sensitive quantum embedding.
Figure 1: Quantum Embedding Model for Word/Sequence Embedding. The model combines: (1) a quantum context-encoding circuit, and (2) a parameterized quantum circuit trained to generate context-sensitive embeddings. Step 1: Define the context {natural, language, is, awesome} and identify the center word (here, processing). The context is transformed into a context matrix. Step 2: Encode the matrix using the context encoding circuit. Context matrices are reshaped for quantum encoding (no resizing required when columns match qubit count). The flexible design adapts to diverse NLP tasks. Step 3: The model predicts the quantum embedding of the center word. Steps 4–6: Compute the loss against the true embedding of the center word.
Figure 2: Architecture of one layer of the quantum context-encoding circuit.
Figure 3: Architecture of one layer of the parameterized quantum circuit.
III-B Model Architecture
This section describes in detail the proposed QCSE model, which leverages parameterized quantum circuits to encode context information and uses its trainable parameters to form a VQC framework. The model consists of two main components: the context-encoding quantum circuit and the ansatz circuit with trainable parameters, as shown in Figure 1. These components are combined to form the full QCSE model, which is used to predict the quantum state for the given word context data.
III-B1 Quantum Context Embedding Circuit
The first component of the QCSE model is the context encoding circuit (shown in blue in Figure 1), which transforms classical data (i.e., the context matrix or vector) into a quantum state. The context encoding circuit, depending on the size of the context matrix, is made up of several layers and each layer is a set of rotation gates with entanglement as shown in Figure 2. Given a context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ (see Sec. IV), we reshape it into $\mathbf{\tilde{C}}\in\mathbb{R}^{2m\times\lceil{n^{2}/2m}\rceil}$. If the number of elements in the last column is less than $2m$, we pad with zeros. Here, $n$ denotes the number of context features and $m$ the number of qubits. The encoding process then applies rotation gates on the qubits to embed the context data into a quantum state, which is subsequently used to learn and predict the corresponding quantum embedding.
To formally describe the encoding process, let $\mathbf{\tilde{C}}=[\mathbf{\tilde{c}}_{1},\mathbf{\tilde{c}}_{2},\dots,\mathbf{\tilde{c}}_{L}]$, where each $\mathbf{\tilde{c}}_{l}\in\mathbb{R}^{2\times m}$ is the set of angles for the quantum gates of one layer applied to the $m$ qubits. The encoding circuit begins by applying a Hadamard gate $H$ to all qubits to create a superposition state, representing the initial uncertainty of the embedding (meaning) of the word. After applying Hadamard gates, the initial state of the quantum system takes the form of:
$\lvert\psi_{0}\rangle=\frac{1}{\sqrt{2^{m}}}\sum_{x\in\{0,1\}^{m}}\lvert x\rangle$ (10)
Subsequently, each vector $\mathbf{\tilde{c}}_{l}$ in the context matrix encodes the rotation angles for the quantum gates, as the $l$-th layer of the encoding circuit. Specifically, for each qubit $q\in\{1,\cdots,m\}$, we apply two rotation gates $RX_{q}(\mathbf{\tilde{c}}_{l}^{(q_{1})})$ and $RZ_{q}(\mathbf{\tilde{c}}_{l}^{(q_{2})})$ around the $x$ and $z$-axis of the Bloch sphere, as follows:
[…]
$=\begin{pmatrix}\cos(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)&-i\sin(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)\\ -i\sin(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)&\cos(\mathbf{\tilde{c}}_{l}^{(q_{1})}/2)\end{pmatrix}$
[…]
$=\begin{pmatrix}\exp(-i\mathbf{\tilde{c}}_{l}^{(q_{2})}/2)&0\\ 0&\exp(i\mathbf{\tilde{c}}_{l}^{(q_{2})}/2)\end{pmatrix}$
where $\mathbf{\tilde{c}}_{l}^{(q_{1})}\text{ and }\mathbf{\tilde{c}}_{l}^{(q_{2})}$ represent the two context parameters for qubit $q$ in each encoding layer. This encoding operation is repeated for each vector $\mathbf{\tilde{c}}_{l}$ in the context matrix, ensuring that the input context is fully encoded into the quantum state.
The full quantum encoding operation for each layer $\mathbf{\tilde{c}}_{l}$ of the context matrix can be described as:
$\lvert\psi_{\text{enc}}\rangle=\prod_{q=1}^{m}\text{$RZ_{q}$}(\mathbf{\tilde{c}}_{l}^{(q_{1})})\text{$RX_{q}$}(\mathbf{\tilde{c}}_{l}^{(q_{2})})\lvert\psi_{0}\rangle,$ (13)
where $\lvert\psi_{0}\rangle$ is the initial quantum state (typically initialized to $\lvert 0\rangle^{\otimes m}$ and then applying $H$-gates as in Eq. 10).
After the single-qubit rotations, the circuit applies a series of controlled-NOT (CNOT) gates between adjacent qubits to introduce entanglement. These CNOT gates are applied in a cascading fashion, connecting qubit $q$ with qubit $q+1$ to generate the entangled quantum state
$\lvert\psi_{\text{ent}}\rangle=\prod_{q=1}^{m-1}\text{CNOT}(q,q+1)\lvert\psi_{\text{enc}}\rangle.$ (14)
This completes the context-encoding process, leaving the quantum system in an entangled quantum state that encodes the input context information to pass throughout the trained parameterized circuit for predicting the quantum context-sensitive embedding of the word.
For illustration purposes, we have added an example in Figure 1. The example illustrates the process of transitioning from a word, processing in this case, and defining its context (natural, language, is, awesome) within the corpus using the bag-of-word approach [60] to define the context of the word (another process can be used, for example, using a character-level process). Then, the context, based on the indexes, positions, frequencies, and distances of words, is transformed into a matrix which is then encoded in a quantum state. This context-sensitive encoding passes through the ansatz for predicting the quantum word embedding, within its context.
III-B2 Ansatz Circuit and Trainable Parameters
The second component of the QCSE model is the ansatz circuit, which comprises trainable parameters. The architecture of the ansatz circuit is shown in Figure 3 in the red block. Let $\mathbf{\Theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\dots,\boldsymbol{\theta}_{M}]$ represent the set of trainable parameters for the ansatz circuit containing $M$ layers. Each layer is composed of unitary operations that depend on the parameters $\boldsymbol{\theta}_{a}$. Each layer parameter set $\boldsymbol{\theta}_{a\in\{1,\cdots,M\}}$ includes two subsets of trainable parameters, $\boldsymbol{\theta}_{a_{1}}\in\mathbb{R}^{2\times m}$ and $\boldsymbol{\theta}_{a_{2}}\in\mathbb{R}^{m-1}$, where $m$ is the number of qubits.
Each ansatz circuit layer consists of two parts: A set of single-qubit rotations, parameterized by $\boldsymbol{\theta}_{a_{1}}$ and a set of controlled-RZ (CRZ) gates, parameterized by $\boldsymbol{\theta}_{a_{2}}$. For each qubit $q$, the ansatz circuit applies the following operations:
$\text{$RX_{q}$}(\theta_{a_{1}}^{(q_{1})})=\exp\left(-i\frac{\theta_{a_{1}}^{(q_{1})}}{2}X_{q}\right),$ (15)
$\text{$RZ_{q}$}(\theta_{a_{1}}^{(q_{2})})=\exp\left(-i\frac{\theta_{a_{1}}^{(q_{2})}}{2}Z_{q}\right)$ (16)
where $\theta_{a_{1}}^{(q_{1})}$ and $\theta_{a_{1}}^{(q_{2})}$ are the two parameters from $\boldsymbol{\mathbf{\theta}_{a_{1}}}$ for the qubit $q$.
Next, the ansatz circuit introduces entanglement using controlled-RZ (CRZ) gates between consecutive qubits, defined as
$\text{CRZ}_{q,q+1}(\phi_{q,q+1})=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&e^{-i\phi_{q,q+1}/2}&0\\ 0&0&0&e^{i\phi_{q,q+1}/2}\end{pmatrix}$ (17)
where $\phi_{q,q+1}$ is the rotation angle parameter from $\boldsymbol{\theta}_{a_{2}}$ applied between qubits $q$ and $q+1$.
The complete ansatz operation for the $a$-th layer can be written as:
$\lvert\psi_{\text{ansatz}}\rangle=\prod_{q=1}^{m-1}\text{$CRZ_{q}$}(\phi_{q,q+1})\prod_{q=1}^{m}\text{$RZ_{q}$}(\theta_{a_{1}}^{(q_{2})})\text{$RX_{q}$}(\theta_{a_{1}}^{(q_{1})})\lvert\psi_{\text{enc}}\rangle$ (18)
This completes the ansatz block, where the parameters $\mathbf{\Theta}$ are trainable and can be optimized to minimize the loss function during the training process.
III-C Combination of the Components
The complete QCSE model combines the context encoding circuit and the ansatz circuit into a single quantum circuit. This quantum circuit is used to perform the final measurement, which predicts the qubit string probabilities for the input data.
After applying both the context-encoding circuit and the ansatz circuit, the expectation value of the Pauli-Z operator is measured on each qubit. For each qubit $q$, the expectation value of the Pauli-$Z$ operator is
$\langle Z_{q}\rangle=\langle\psi_{\text{ansatz}}\rvert Z_{q}\lvert\psi_{\text{ansatz}}\rangle.$ (19)
This quantity takes values in the interval [1,1][-1,1]. The measurement probabilities of $q$, in the computational basis states, are
$P(\lvert 0\rangle_{q})=\frac{1+\langle Z_{q}\rangle}{2},\quad P(\lvert 1\rangle_{q})=\frac{1-\langle Z_{q}\rangle}{2}.$ (20)
Finally, the output of the quantum embedding model is the list of probabilities $P(\lvert 1\rangle_{q})$ (or equivalently $P(\lvert 0\rangle_{q})$) for each qubit $q$, which represent the model’s prediction for the encoded context.
The quantum embedding model described in this section provides a novel method for encoding classical data into quantum states using parameterized quantum circuits. The combination of context encoding and trainable ansatz layers allows for a flexible and expressive quantum model that can be optimized to minimize a given loss function, such as the cross-entropy. The output of the model is a set of bitstring probabilities, which can be used in various quantum machine learning applications. Here, we consider it in the context of QNLP. Figure 1 visualizes the complete model and the whole training process. The process establishes a basic context by leveraging nearby words or characters to generate a context matrix that encodes the quantum state of the word for further processing.
III-D Complexity Analysis of the Model
In QCSE, the number of qubits required for the quantum encoding process is driven primarily by the size of the vocabulary, not by the context size. This is because the context matrix, rather than increasing the number of qubits, is flattened, reshaped and padded with zeros when required, and encoded sequentially across multiple layers of the quantum circuit. The same set of qubits is reused for the encoding, ensuring that the qubit requirement remains constant.
The number of qubits, $m$, required by the model is governed by the size of the vocabulary $|V|$. If each word in the vocabulary is represented by a computational basis state, the number of qubits is proportional to the logarithm of the vocabulary size,
$m=\lceil\log_{2}|V|\rceil.$
For a large vocabulary of size $|V|\sim 10^{5}$, such as the english vocabulary, this results in $m\sim 17$ qubits.
While the number of qubits remains constant, the quantum gate complexity increases with the size of the context window and the desired depth of the ansatz layers. The complexity of the quantum gate operations can be broken down into two primary components: (1) the context encoding block and (2) the ansatz block.
  • Context Encoding Block Complexity: Each layer corresponds to one portion of the context matrix being mapped onto the qubit system. Since the number of qubits is fixed, the encoding of the context matrix is distributed across multiple layers. Therefore, the complexity of the gates of the context encoding block scales linearly with the number of context encoding layers $L$. The total number of gates for the context encoding block is
    $G_{\text{context}}=(3m-1)\cdot L.$ (21)
  • Ansatz Block Complexity: The ansatz block is repeated $n_{\text{repeats}}$ times. For a system of $m$ qubits, each layer of the ansatz consists of $2m$ single-qubit rotation gates and m1m-1 controlled entangling gates (such as CNOT or CRZ). Thus, the number of gates in a single ansatz layer is approximately $G_{\text{layer}}=3m-1$. The total number of gates required for the ansatz block with $n_{\text{repeats}}$ repetitions is
    $G_{\text{ansatz}}=(3m-1)\cdot M.$ (22)
The total gate complexity of the entire model is the sum of the gates required for the ansatz block and the context encoding block, i.e,
$G_{\text{total}}=G_{\text{ansatz}}+G_{\text{context}}=(3m-1)(M+L)+m.$ (23)
IV Computing the Context Matrices and Vectors
This section details approaches to generate unique context matrices. These methods are grounded in various transformations, such as the hash function [61], exponential decay, and sinusoidal functions[62, 63]. The aim of these functions is to encode word relationships within context windows, producing matrices that capture word identities, distance and positional information, and semantic-syntactic features. Then, these matrices are reshaped to $\mathbf{\tilde{C}}\in\mathbb{R}^{2m\times\lceil{n^{2}/2m}\rceil}$ and padded with zeros to match the layers.
IV-A Exponential Decay with Sinusoidal Encoding
This method defines the context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ for a context window of $n$ words by incorporating both positional distance and semantic relationships through sinusoidal transformations. For words at positions $i$ and $j$ in the context window, each element $c_{ij}$ is computed as
$c_{ij}=e^{-\alpha|i-j|}\sin(\omega\theta_{i})\cos(\omega\theta_{j})+\theta_{i}$ (24)
where:
  • $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ is the normalized angle for word at position $i$.
  • $\text{idx}_{i}$ is the vocabulary index of the word at position $i$.
  • |ij||i-j| represents the positional distance between words.
  • $\alpha>0$ controls the exponential decay rate with distance.
  • $\omega>0$ determines the frequency of sinusoidal components.
  • $|V|$ is the vocabulary size or maximum index value.
This formulation combines three key components: (1) an exponential decay term that captures the positional proximity between words, (2) a sinusoidal encoding that provides distinct representations for different words, and (3) an index-based uniqueness term that ensures different words receive distinct representations even when they appear at the same relative positions.
IV-B Index-Based Diagonal Encoding
The second approach introduces a specialized treatment for diagonal elements to emphasize self-relationships while maintaining positional decay for off-diagonal elements. The context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ is defined as
$c_{ij}=\begin{cases}\log(1+\text{idx}_{i}),&\text{if }i=j\\ e^{-\alpha|i-j|}\sin(\omega\theta_{i})+\theta_{i},&\text{if }i\neq j\end{cases}$ (25)
where:
  • $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ is the normalized angle
  • $\log(1+\text{idx}_{i})$ provides a logarithmically scaled value for diagonal element and assuming that indexes start from 0.
  • Other parameters follow the definitions from the previous method.
The logarithmic scaling of diagonal elements ensures that words with higher indices (presumably unknown or less common words) receive stronger self-relationships, which can be valuable for preserving the significance of rare terms in the context representation.
IV-C Positional Phase Shift Encoding
This method introduces position-dependent phase shifts in the sinusoidal components, creating a rich representation that encodes both absolute positions and relative distances. The context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ is defined as
$c_{ij}=e^{-\alpha|i-j|}\sin(\omega i+\delta\theta_{i})+\theta_{i}$ (26)
where:
  • $\delta>0$ is a phase shift scaling parameter.
  • $\omega i$ introduces a position-dependent base phase.
  • $\delta\theta_{i}$ adds a word-specific phase shift.
  • Other parameters follow previous definitions.
This formulation effectively creates a unique encoding for each word based on both its identity and its position in the context window, enabling the model to distinguish between identical words appearing at different positions.
IV-D Hash-Based Modulation Encoding
To enhance the uniqueness of representations and reduce potential collisions between words with similar indices, this method introduces a hash-based modulation approach. The context matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ is defined as
$c_{ij}=e^{-\alpha|i-j|}\sin(\omega i+h_{i})+\theta_{i}$ (27)
where:
  • $h_{i}=(\text{idx}_{i}\cdot p)\bmod N$ is a simple hash function
  • $p$ is a prime number
  • NN is the hash space size
  • Other parameters follow previous definitions
The hash-based modulation introduces an additional layer of uniqueness that is particularly valuable for handling large vocabularies where index-based collisions might otherwise occur. The use of prime numbers in the hash function helps ensure a good distribution of hash values across the vocabulary.
IV-E Positional Angular Shift Vector Representation
While the previous methods focused on matrix representations, a context-to-vector method introduces a vector-based context representation that efficiently captures word relationships. A context window of $n$ words, is defined as context vector $\mathbf{v}\in\mathbb{R}^{n}$ where each element $v_{i}$ is
$v_{i}=\omega\theta_{i}$ (28)
where:
  • $\theta_{i}=\text{idx}_{i}\cdot\frac{2\pi}{|V|}$ is the normalized angle
  • $\omega>0$ determines the frequency of angular encoding
  • $\text{idx}_{i}$ is the vocabulary index of the word at position $i$
This vector representation provides a more compact and straightforward encoding that emphasizes word identities through angular transformations while incorporating positional information implicitly through the qubits ordering or indices.
Each of these methods offers different trade-offs between computational complexity, representational capacity, and numerical stability.
V Evaluation of QCSE
The QCSE model was implemented and tested using PennyLane [^1] https://pennylane.ai/ in two separate evaluations. The first (Section V-A) assessed five context encoding methods (Section IV) to identify the most effective approach. The second (Section V-B) compared the quantum embedding model with word2vec’s CBOW [64] to examine potential quantum advantages in QCSE.
V-A Evaluation of Context Encoding Methods
This section outlines the evaluation setup and results for assessing the QCSE model’s performance with different context encoding methods across two corpora.
V-A1 Evaluation Setup
The evaluations were structured as follows. Two datasets were used:
  • English: a vocabulary size of $31$ and 656 center-context word pairs.
  • Fulani (a low-resource African language) [65]: a vocabulary size of $27$ and 128 word-context pairs.
The model used 5 qubits, with each word encoded as a quantum state. The quantum ansatz had 3 layers (14 trainable parameters each, totaling 42 parameters). A context window of 2 (4-word span) was applied, following the method in Figure 1.
V-A2 Results and Discussion
(a) English dataset (656 pairs)
(b) Fulani dataset (128 pairs)
Figure 4: Training loss across methods: (a) Exponential Decay Sinusoidal, (b) Index-based Diagonal, (c) Positional Phase Shift, (d) Hash-based, (e) Positional Angular Shift. Fulani represents a low-resource African language.
The training process optimizes embedding similarity via cross-entropy loss minimization, and we compare convergence rates to find the method with the fastest stable loss reduction. The results, Figure 4(a) and Figure 4(b) for English and Fulani datasets respectively, demonstrate that Exponential Decay Sinusoidal and the Positional Phase Shift encoding methods consistently yield the best performance across both datasets. In the English dataset, the Hash-based and Index-based Diagonal methods achieve the lowest final training loss, around $4.87$. The Positional Angular Shift method performs worst, fluctuating around $5.2$, indicating poorer learning of context. These trends suggest that smooth, structured encodings—such as sinusoidal or phase-based—enable more effective learning of contextual relationships.
In the Fulani dataset (Figure 4(b)), where only 128 word-context pairs were used. The Exponential Decay Sinusoidal encoding reaches a final loss of about $6.1$, the lowest among all methods, while the Hash-based method stabilizes around $6.2$. Other methods like Index-based Diagonal and Hash-based settle around $6.3$, while Positional Angular Shift remains the worst performer at approximately $6.6$ throughout training. The sharper performance gap in the low-resource setting highlights the robustness of the Exponential Decay Sinusoidal encoding, as it can still effectively represent positional context even with limited data. This could support its suitability for both high- and low-resource NLP applications within quantum embedding models.
The success of these methods can be attributed to their inherent compatibility with quantum systems. Exponential Decay Sinusoidal leverages periodicity and decay to encode positional information. This performance underscores the importance of encoding methods that balance quantum native operations with classical NLP principles (e.g., positional awareness), especially when scaling across dataset sizes. The next section analyses the performance of the QCSE model using the exponential decay sinusoidal method for context encoding.
V-B Analysis of QCSE model performance
This section presents a comprehensive analysis of QCSE model performance compared to classical word embedding approaches, specifically the well-known CBOW model[64]. The evaluation focuses on parameter efficiency, scalability characteristics.
V-B1 Evaluation Setup
Evaluations were carried out on a data set of vocabulary size 34, comprising 300 sentences (1200 words total), from which 1200 center-context word pairs were extracted. The data set was systematically partitioned into training and testing subsets to ensure robust evaluation. The QCSE model architecture employs 6 qubits with varying layer depths (1-8 layers), while the classical CBOW baseline utilizes vector dimensions of 20 and 50. All models were evaluated using similarity metrics on word relationship tasks, providing a standardized measure of semantic representation quality.
V-B2 Results and Discussion
The dataset is split into training and test sets. The model is trained to maximize similarity between true and predicted embeddings via cross-entropy loss. During validation, accuracy is measured using the hamming distance to compute the similarity scores by checking if the predicted word or embedding matches the center of the context. The predictions were considered accurate if the similarity between quantum embeddings exceeded $0.90$, equivalent to a Hamming distance of at most 1 qubit difference. Table I shows that the QCSE model exhibits distinctive performance characteristics as a function of the depth of the circuit. In this evaluation, the optimal architecture is identified with 6 layers, achieving 37.82% accuracy with only 102 trainable parameters and outperforms the classical model (37.82% vs 34.78%). This represents a clear performance peak, with both shallower and deeper architectures showing reduced effectiveness, as reflected in the curves. The performance curve reveals several critical insights into quantum word embedding behavior: (1) Shallow Circuit Limitations: Single-layer QCSE (9.58% accuracy) demonstrates insufficient expressibility for capturing complex semantic relationships. The rapid improvement from 1 to 4 layers (9.58% to 36.67%) indicates that quantum circuits require adequate depth to develop meaningful word representations. (2) Optimal Depth Identification: The performance plateau between 4-6 layers, with peak performance at 6 layers (37.82%), suggests an optimal balance between circuit expressibility and trainability. This finding is crucial for practical quantum NLP applications, as it identifies the minimal circuit depth required for competitive performance with respect to the data set size. (3) Overparameterization Effects: Performance degradation beyond 6 layers (37.82% to 34.58% at 8 layers) indicates quantum circuit over-parameterization, analogous to classical deep learning phenomena. This suggests that quantum models are subject to similar optimization challenges as their classical counterparts, requiring careful architectural design. The results of the classical CBOW model reveal interesting scalability characteristics that further emphasize the quantum parameter efficiency. Counter-intuitively, increasing the CBOW vector dimension from 20 to 50 (increasing parameters from 680 to 1700) with the same epochs resulted in decreased performance from 34.78% to 20.00%. This performance degradation suggests that, just like the quantum model, classical models suffer from optimization difficulties when scaled inappropriately for limited training data[64].
These results demonstrate the parameter utilization efficiency of the QCSE model compared to the CBOW model. The optimal QCSE model (6 layers, 102 parameters) outperforms the classical CBOW model with 680 parameters, achieving 3.04% higher accuracy with 85% fewer parameters. This represents a parameter efficiency improvement when measured by accuracy per parameter. When comparing against the larger CBOW configuration (50-dimensional, 1700 parameters), the quantum model achieves 17.82% higher accuracy (37.82% vs. 20.00%) while using 94% fewer parameters. This contrast highlights the ability of the quantum model to capture semantic relationships with fewer parameters. The parameter efficiency becomes particularly relevant in resource-constrained environments where parameter count directly impacts computational overhead and memory requirements.
TABLE I: Comparative Performance Analysis of Pretrained QCSE and Classical Continuous Bag of Words (CBOW) Models
Model Number of Number of Number of Number of Accuracy
layers qubits/Vec trainable Training [%]
tor dimension params words (similarity
score)
QCSE 1 6 17 1200 9.58
2 6 34 1200 27.09
3 6 51 1200 32.50
4 6 68 1200 36.67
5 6 85 1200 32.50
6 6 102 1200 37.82
7 6 119 1200 36.67
8 6 136 1200 34.58
CBOW 20 680 1200 34.78
50 1700 1200 20.00
The training loss curves presented in Figure 5 provide insights into the learning dynamics, convergence behavior, and optimization characteristics of both quantum and classical word embedding models. The QCSE models exhibit different training behaviors as a function of circuit depth, with each layer configuration demonstrating unique convergence characteristics that correlate with final performance outcomes.
The single-layer quantum circuit displays unstable training dynamics with oscillatory behavior throughout the training process. The loss remains constant around 8.8, never achieving convergence. This behavior directly corresponds to the poor final accuracy of 9.58%, indicating that single-layer quantum circuits lack sufficient expressibility to capture meaningful semantic relationships. As circuit depth increases from 2 to 4 layers, we observe a progressive improvement in training stability and convergence behavior. The 2-layer circuit shows rapid initial convergence from approximately 6.4 to 6.17 within the first 10 epochs, followed by stable plateauing around 6.2. This convergence pattern correlates with the performance improvement to 27.09% accuracy. The 3-layer configuration demonstrates stable training dynamics, with a loss reduction from approximately 6.0 to 5.55 and achieving 32.5% accuracy. The 4-layer model exhibits stable training trajectory among shallow circuits, with consistent loss reduction from about 5.9 to 5.55 and minimal oscillations, corresponding to 36.67% accuracy.
The 6-layer quantum circuit, which achieved the highest accuracy of 37.82%, displays a best training dynamics characterized by monotonic loss reduction from approximately 6.0 to 5.5. The convergence curve shows three distinct phases: rapid initial descent (epochs 0-20). Deeper quantum circuits (7-8 layers) exhibit similar final loss values but with slightly reduced accuracy.
The 20-dimensional CBOW model (680 parameters and 34.78% accuracy) exhibits a training dynamics with fluctuations throughout the training process. The loss fluctuates between approximately 6.9 and 6.35, showing periodic variations. The higher-dimensional CBOW model (1700 parameters) shows similar training oscillations but achieves significantly lower performance (20.00% accuracy). The training curve displays persistent oscillations with loss values fluctuating around 6.4, indicating that increased model dimensionality does not translate to improved optimization or performance for this dataset size[64].
Compared to classical models such as word2vec[64], GloVe[44, 32], which typically require high-dimensional vectors and large data sets, the proposed quantum model demonstrates the ability to perform well on large and small datasets. Furthermore, compared to other quantum embedding models that rely on classical embeddings as input[47, 32], QCSE model learns directly from scratch. This significantly reduces the dependence on classical pre-trained embeddings and highlights the efficiency of using quantum language modeling.
This suggests that quantum parameterized circuits, when properly designed, may possess more favorable optimization geometries than their classical counterparts. Our quantum model can achieve stable convergence with significantly fewer parameters. The proposed quantum circuits require fewer epochs to reach stable convergence, with most QCSE models achieving plateau performance within 10-30 epochs compared to the continued instability of classical models throughout 50 epochs.
Figure 5: Training loss curves that analyzes both QCSE and CBOW behaviors.
VI Applications and Future Directions
The proposed quantum embedding model enables key NLP applications , such as language modeling, translation, and text generation by capturing linguistic relationships. QCSE can be extended to quantum encoder-decoder frameworks [66, 67, 68, 69] for tasks like translation and summarization. Quantum states encode context-sensitive embeddings, while decoders reconstruct classical outputs, leveraging quantum parallelism for efficiency. Quantum self-attention [68, 69] could compute superposition-based attention scores, reducing the quadratic complexity of classical transformers [63]. Integrating quantum embeddings into pretrained models (e.g., BERT, GPT [70]) may enhance context sensitivity.
Beyond bag-of-words, Skip-Gram [60] or dependency-based contexts could enrich training. Attention mechanisms [63] may dynamically weight context words, enabling quantum transformers with masked/causal language modeling. QCSE benefits low-resource languages [71, 72] (e.g., Fulani) by extracting richer semantics from limited data.
Key directions include:
  • Impact of Circuit Depth: Increasing the number of layers in the word-context encoding circuit or the ansatz could enhance the model’s capacity to capture more complex semantic structures. A systematic study of different circuit depths would help identify the trade-offs between circuit expressiveness and noise-induced errors.
  • Different Ansatz Exploration: Different types of parameterized quantum circuits [73] should be explored. A more suitable ansatz could potentially improve the encoding of intricate word relationships, especially in low-resource settings.
  • Train the QCSE model on large-scale datasets using varied training parameters to evaluate its scalability, performance, and robustness across different configurations.
VII Conclusion
This work presented a pretrained Quantum Context-Sensitive Embedding (QCSE) model that leverages quantum computing principles to address NLP challenges. The study demonstrated the feasibility of applying quantum circuits to word embedding tasks, showing strong potential for capturing semantic relationships that classical models may struggle with. The approach encodes word–context relationships as quantum states, exploiting quantum properties such as superposition and entanglement to represent linguistic patterns. Evaluations on both medium- and small-sized datasets show that the model generalizes well, with notable success in low-resource language scenarios. This suggests that quantum models can mitigate the data scarcity problem often faced in NLP, learning meaningful semantic relationships even from limited data. Overall, this work establishes a foundation for further exploration of quantum word embeddings, particularly for low-resource NLP applications.
Acknowledgements
Charles M. Varmantchaonala is grateful for financial support via a research grant from the German Academic Exchange Service (DAAD).